OpenStudy (anonymous):

A diver is on a board 1.80 m above the water. She jumps straight up at 3.62 m/s. What is her maximum height above the water?

OpenStudy (anonymous):

The Kinetic energy at the moment she jumps is 1/2 mv^2 .. right ?

OpenStudy (anonymous):

this is the equation for kinetic energy, but this is an energy problem

OpenStudy (anonymous):

Wait man.. The potential energy of diver in air at maximum height = mgh (h = max.height)

OpenStudy (anonymous):

By conservation of mechanical energy, you have mgh = 1/2 mv^2 h = v^2/2g = (3.62^2)/(2*9.8) = 0.67 m This is height above the board Height above water = 0.67 + 1.80 = 2.47 m Got it ?? @DejohnetteW

OpenStudy (anonymous):

Yes! thank you so much!

OpenStudy (anonymous):

Welcome bro !

OpenStudy (anonymous):

can you help me again? A stone is dropped from a height of 7.70m. How fast is it going when it reaches the ground?